# Stellar-mass black holes in young massive and open stellar clusters – VI. Role of external galactic field

### v1

@@ Acknowledgements.
The author (SB) acknowledges funding for this work by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the project “The dynamics of stellar-mass black holes in dense stellar systems and their role in gravitational wave generation” (project number 405620641; PI: S. Banerjee). All $N$-body simulations have been carried out on the gpudyn-series GPU computing servers containing NVIDIA Ampere A40 and NVIDIA RTX 2080 GPUs, located at the AIfA, University of Bonn. The gpudyn servers have been sponsored by the above-mentioned DFG project and the HISKP. SB is thankful to the IT teams of the AIfA, HISKP, and University of Bonn HPC center for their generous support.

### v2

@@ Acknowledgements.
The author (SB) thanks the reviewer for their constructive criticisms and suggestions that have helped to improve the paper. SB acknowledges funding for this work by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the project “The dynamics of stellar-mass black holes in dense stellar systems and their role in gravitational wave generation” (project number 405620641; PI: S. Banerjee). All $N$-body simulations have been carried out on the gpudyn-series GPU computing servers containing NVIDIA Ampere A40 and NVIDIA RTX 2080 GPUs, located at the AIfA, University of Bonn. The gpudyn servers have been sponsored by the above-mentioned DFG project and the HISKP. SB is thankful to the IT teams of the AIfA, HISKP, and University of Bonn HPC center for their generous support.